OpenStudy (kva1992):

sqrt{x-2} + 3 = x how do you solve that? answer is \frac{ 7 pm sqrt{5} }{ 2 }

OpenStudy (kva1992):

\[\frac{ 7\pm \sqrt{5} }{ 2 }\] thats the answer

OpenStudy (kva1992):

i ended up with 77 in the square root any idea what i did wrong?

OpenStudy (kva1992):

did that

OpenStudy (mathmath333):

\(\large\tt \color{black}{\sqrt{x-2} = x -3}\) \(\large\tt \color{black}{x-2 = (x -3)^2}\)

OpenStudy (kva1992):

youd end up with x^2 -7x-7

OpenStudy (kva1992):

use the quadratic formula and im still getting the same as before

OpenStudy (kva1992):

cuz its 49 + 28 and that adds up too 77

OpenStudy (mathmath333):

\(\large\tt \color{black}{it~~will~~be~~x^2 -7x+11}\)

OpenStudy (kva1992):

x-3 squared is -9

OpenStudy (kva1992):

you cant have a positive

OpenStudy (kva1992):

unless im missing a step

OpenStudy (mathmath333):

\(\large\tt \color{black}{(a-b)^2=a^2-2ab+b^2}\)

OpenStudy (kva1992):

yeah its x^2 - 6x -9 its the property of squares

OpenStudy (kva1992):

wait really?

OpenStudy (kva1992):

damn im just droppen the ball

OpenStudy (mathmath333):

\(\large\tt \color{black}{there~~will~~(+)~~sign~~before~~9}\)

OpenStudy (kva1992):

well that solves my problem lol i can figure it out now haha thanks for the help lol

OpenStudy (mathmath333):

\(\large\tt \color{black}{x^2 - 6x +9}\) welcome